OpenStudy (anonymous):

What equation represents the line passing through the point (-5, 1) with a slope of 3/2.

OpenStudy (itsonlycdeee):

\[y =\frac{ 3 }{ 2 }x +b \] \[1 = \frac{ 3 }{ 2 }* (-5) +b \rightarrow 1 = \frac{ -15 }{ 2 } + b \] b = 17/ 2, therefore...?